Your firm has identified three potential investment projects. The projects and their cash flows are shown​ here:  ​(Click on the icon located on the​ top-right corner of the data table below in order to copy its contents into a​ spreadsheet.)

Project

Cash Flow Today

​(millions)

Cash Flow in One Year

​(millions)

A

-$15

$$16

B

$4

$3

C

$17

−$11

Suppose all cash flows are certain and the​ risk-free interest rate is

6%.

a. What is the NPV of each​ project?

b. If the firm can choose only one of these​ projects, which should it​ choose?

c. If the firm can choose any two of these​ projects, which should it​ choose?

a. What is the NPV of each​ project?

The NPV of project A is

million. ​(Round to two decimal​ places.)

The NPV of project B is

million.  ​(Round to two decimal​ places.)

The NPV of project C is

million. ​(Round to two decimal​ places.)

b. If the firm can choose only one of these​ projects, which should it choose based on the NPV decision​ rule?  ​(Select the best choice​ below.)

A. Project C

B. Project B

C. Project A

D. Cannot Tell

c. If the firm can choose any two of these​ projects, which should it choose based on the NPV decision​ rule?  ​(Select the best choice​ below.)

A.Project B and project A

B.Project A and project C

C.Project B and project C

D.

Cannot tell
Answer #1

NPV = present value of cash inflows – present value of cash outflows

Project A = 16/(1.06) – 15

= $0.09434 millions

Project B = 3/1.06 + 4

= 6.83 million

Project C = -11/1.06 + 17

= 6.62 million

B.Project B as it has highest NPV

C.Project B and C